Accelerator Update - July 30-August 1
Pbar stopped stacking due to a controller problem at 8:20 AM. Operators contacted an expert and then prepared the tunnel for access. The expert replaced a power supply for a Programmable Logic Controller at 9:22 AM.
ComEd admitted that they had suffered two power glitches this morning at 3:18 AM and 5:19 AM. This accounts for the two Booster power supply trips (WAPS) this morning.
MI expert began a study period at 8:55 AM.
While the Linac and Pbar work was occurring, another Pbar expert repaired a Pbar Debuncher (DRF1) at 9:41 AM.
Linac beam resumed at 9:44 AM. MI and MiniBooNE beam resumed at 9:52 AM. Stacking resumed at 10:23 AM.
Earlier, a Booster RF station (BRF11) was brought back online. It tripped off twice at 10:47 AM and then at 10:49 AM and once again at 11:47 AM.
A Pbar Debuncher power supply (D:H301) tripped off at 11:26 AM. An expert replaced the supply and had it back in operation 35 minutes later.
At 1:30 PM, Operations received a fire alarm for the Tevatron (TeV) sector F0 service building. At the same time a F0 RF power supply caused a trip that aborted store 2830. RF expert investigated but could find nothing wrong. The thought was that maybe a two-way radio had caused the abort, but tests with the Fire Department didn't prove that out.
Operations put the TeV through a dry squeeze at 2:35 PM. TeV experts then began a study period at 3:07 PM.
Thursday July 31
D0 came out of access 12:02 AM.
Vacuum experts replace two controller cards at 12:31 AM.
Operations put the TeV through a dry squeeze at 2:05 AM.
Two Booster RF stations (BRF5 & 6) tripped off due to a vacuum burst. Everything reset okay.
A TeV study period ran from 3:30 AM to 5:51 AM and then went into shot setup.
At 8:10 AM, cryo technicians confirmed a valve problem with the wet engine at the sector A4 service building. They had the valve adjustment completed by 8:31 AM.
The TeV suffered a sector A1 quench at 8:57 AM. This occurred while ramping up for the next store. The store was lost. Investigators found no reason for this trip.
Pbar resumed stacking at 9:19 AM.
A MI vacuum burst occurred at the NuMI stub at 5:54 PM. This halted stacking until 6:15 PM.
At 7:07 PM, the TeV suffered a sector F1 quench. A console in the Main Control Room froze at the beginning of a load proton aggregate.
A lightening and rainstorm rushed across the lab around 10 PM, which caused a few trips, alarms, and havoc for the operators, but nothing major happened.
Friday August 1
Operations put the TeV into shot setup and at 3:52 AM established store 2847 with an initial luminosity of 3.5E31. Pbar resumed stacking.
CDF had a problem at 3:57 AM with equipment that lets them know when it's okay to start taking data. This problem was resolved by 5:46 AM.
At 5:32 AM, Booster beam problems required Operations to lessen the MiniBooNE beam intensity. This was still the case at 8 AM.
Comments and Suggestions
|last modified 8/5/2003 email Fermilab|